# Analytical and Numerical Studies of Dark Current in Radiofrequency Structures for Short-Pulse High-Gradient Acceleration

### v1

@@ II Theory of Electron Multipacting
In this work, we study electron multipacting in crossed RF fields near the sidewall of a pillbox cavity, focusing on closed electron trajectories that originate from and return to the sidewall. These trajectories are of particular interest, as they are resonant with the RF field and are expected to play a dominant role in the multipactor phenomenon. Our analysis includes an axial electric field $E_{z}$, a radial electric field $E_{r}$, and an azimuthal magnetic field $B_{\theta}$. The radial component $E_{r}$ arises near the cavity sidewall due to the presence of the cavity iris. We modify the field distribution in the pillbox cavity to approximate conditions near the sidewall region of the $X$-band photogun cavities. We assume an axisymmetric field distribution, so electron motion is confined to the $r$$z$ plane. This assumption holds as long as any asymmetries or perturbations that could induce out-of-plane motion remain negligible. These approximations yield predictions in good agreement with simulations of the actual cavity, as presented in Sec. III. We begin by calculating secondary electron trajectories and then evaluate the SEY in both single-impact and multipacting resonance conditions.
@@ II.1 Secondary electron trajectory calculations
Figure 4: Comparison of RF field amplitudes from cst simulations and analytical approximations for the full-cell cavity with radius $a=9.49$ mm and length $L=6.50$ mm. The field amplitudes—$B_{\theta}$ in (a), $E_{r}$ in (b), and $E_{z}$ in (c)—are all scaled to an axial field amplitude of $E_{z0}=100$ MV/m at the cavity center. The shaded region in (c) indicates the radial range near the sidewall where the analytical field approximation is applied. Solid red curves represent cst simulation data, while black dashed lines indicate the analytical field amplitudes used in the model.
@@
where $E_{z0}$ again denotes the $E_{z}$ field amplitude at the cavity center, and $P=(\mathrm{d}E_{z}/\mathrm{d}r)/E_{z0}$ The azimuthal magnetic field $B_{\theta}$ is modeled as:
@@ II.2 Single-cycle SEY calculations
To compute the SEY, we begin by extracting the impact energy $W_{i}$ and incidence angle $\theta_{i}$ from the electron trajectory. The axial impact location is denoted by $z_{i}$, and the RF phase at impact is given by $\phi_{i}=\omega\tau+\phi_{0}$, where $\tau$ is the time of flight. We assume an initial kinetic energy of 2 eV for secondary electrons, consistent with the emission model in Ref. [30], which corresponds to the minimum energy needed to overcome the initial retarding field. We then perform a parameter sweep over the electron emission angle $\alpha_{0}$ and initial RF phase $\phi_{0}$, across a range of axial emission positions $z_{0}$ and field amplitudes characterized by the axial electric field amplitude at the cavity center, $E_{z0}$. All field components—$E_{z}$, $E_{r}$, and $B_{\theta}$—are scaled proportionally with $E_{z0}$, which serves as the reference parameter for the sweep. SEY values exceeding unity are retained and averaged to generate a representative single-cycle SEY contour map. The resulting map is shown in Fig. 5, highlighting regions in $E_{z0}$$z_{0}$ space that are susceptible to enhanced secondary emission and the potential onset of electron multipacting.
@@ III.1 Field Emission Simulations
Figure 9: Multipacting simulation setup at the middle cutting plane of the $X$-band cavities, divided into the full cell, and Region I and Region II in the half cell. The reference points $P_{1}$, $P_{2}$ and $P_{3}$ denote the centers of each cavity, respectively. The relevant dimensions are: $R$1 = 8.09 mm, $R$2 = $R$3 = 9.49 mm, $Z$1 = 2.40 mm, $Z$2 = 2.60 mm, and $Z$3 = 6.50 mm.
[…]
@@ III.2 Multipacting simulations in the short-pulse regime
[…]
@@ Appendix A Simplified theory of multipactor in crossed RF fields with closed-form solutions
Table 1: Comparison of $E_{z0}$ ranges for various multipacting resonance orders $N$, calculated using the simplified and full analytical models. The simplified model captures the scaling trend well, particularly for higher-order resonances ($N\geq 2$).

### v2

@@ II Theory of Electron Multipacting
In this work, we study electron multipacting in crossed RF fields near the sidewall of a pillbox cavity, focusing on closed electron trajectories that originate from and return to the sidewall. These trajectories are of particular interest, as they are resonant with the RF field and are expected to play a dominant role in the multipactor phenomenon. Our analysis includes an axial electric field $E_{z}$, a radial electric field $E_{r}$, and an azimuthal magnetic field $B_{\theta}$. The radial component $E_{r}$ arises near the cavity sidewall due to the presence of the cavity iris. We modify the field distribution in the pillbox cavity to approximate conditions near the sidewall region of the $X$-band photogun cavities. We assume an axisymmetric field distribution, so electron motion is confined to the rrzz plane. This assumption holds as long as any asymmetries or perturbations that could induce out-of-plane motion remain negligible. These approximations yield predictions in good agreement with simulations of the actual cavity, as presented in Sec. III. We begin by calculating secondary electron trajectories and then evaluate the SEY in both single-impact and multipacting resonance conditions.
@@ II.1 Secondary electron trajectory calculations
[…]
@@
where $E_{z0}$ again denotes the $E_{z}$ field amplitude at the cavity center, and $P=(\mathrm{d}E_{z}/\mathrm{d}r)/E_{z0}$ . The azimuthal magnetic field $B_{\theta}$ is modeled as:
@@ II.2 Single-cycle SEY calculations
Figure 6: Sample multipacting resonance modes ($N=1\text{--}7$) calculated for the RF field distribution in the full cell of the $X$-band photocathode cavities. The corresponding axial field amplitudes $E_{z0}$ at the cavity center for each order are 106, 61, 39, 31, 26, 22, and 19 MV/m, respectively.
@@ III.1 Field Emission Simulations
Figure 9: Multipacting simulation setup at the middle cutting plane of the $X$-band cavities, divided into the full cell, and Region I and Region II in the half cell. The reference points $P_{1}$, $P_{2}$ and $P_{3}$ denote the centers of each cavity, respectively. The relevant dimensions are: RR1 = 8.09 mm, RR2 = RR3 = 9.49 mm, ZZ1 = 2.40 mm, ZZ2 = 2.60 mm, and ZZ3 = 6.50 mm.
Figure 10: Multipacting resonance trajectories simulated in the full cell (defined in Fig. 9). The corresponding axial field amplitudes $E_{z0}$ for resonance orders $N=1\text{--}7$, evaluated at point $P_{3}$ (the center of the full cell), are 106, 61, 39, 31, 26, 22, and 19 MV/m, respectively. These trajectories show good agreement with the analytical results in Fig. 6, as the RF field distribution in the full cell is well approximated by the analytical model.
Figure 11: Multipacting resonance trajectories in the half-cell for (a) Region I and (b) Region II. In Region I, the axial fields $E_{z0}$ evaluated at point $P_{1}$ for resonance orders $N=1\text{--}4$ are 112, 61, 39, and 28 MV/m, respectively. In Region II, the axial fields $E_{z0}$ evaluated at point $P_{2}$ for resonance orders $N=1\text{--}5$ are 95, 46, 30, 25, and 21 MV/m, respectively.
@@ III.2 Multipacting simulations in the short-pulse regime
Figure 12: Temporal evolution of the electric field amplitudes and the number of multipacting electrons for (a) Region I, (b) Region II, and (c) the full cell. In each subplot, the on-axis electric field at point $P_{1}$, $P_{2}$, or $P_{3}$ (red solid line), the electric field across the photocathode surface (black dashed line), and the number of multipacting electrons (blue solid line) are shown. For Region I, the dark current peaks corresponding to $N=1\text{--}4$ appear at 113, 66, 43, and 30 MV/m; for Region II, the peaks for $N=1,2,4,5$ are at 100, 47, 28, and 20 MV/m; and for the full cell, the peaks for $N=1,2,3,5$ are at 113, 51, 35, and 27 MV/m. Higher-order resonance modes ($N\geq 6$) also contribute to later peaks in the full cell.
@@ Appendix A Simplified theory of multipactor in crossed RF fields with closed-form solutions
Table 1: Comparison of $E_{z0}$ ranges for various multipacting resonance orders NN, calculated using the simplified and full analytical models. The simplified model captures the scaling trend well, particularly for higher-order resonances ($N\geq 2$).